Three parallel connected resistors when connected across a d.c. voltage source dissipate a total power of 72 W. The total current drawn is 6 A, the current flowing through the first resistor is 3 A and the second and third resistors have equal value. What are the resistances of the three resistorsThree parallel connected resistors when connected across a d.c. voltage source dissipate a total power of 72 W. The total current drawn is 6 A, the current flowing through the first resistor is 3 A and the second and third resistors have equal value. What are the resistances of the three resistorsThree parallel connected resistors when connected across a d.c. voltage source dissipate a total power of 72 W. The total current drawn is 6 A, the current flowing through the first resistor is 3 A and the second and third resistors have equal value. What are the resistances of the three resistors

Grade:12th pass

1 Answers

Khimraj
6 years ago
Let the total resistance is Req and dc voltage is V voltage and total current be I amp then
Total power P = 72W and I = 6A
V*I = 72
THen V = 12V
Current through first resistor is I1 3A let it resistance is R1 then
12/R1 = 3
R1= 4
Let second resistance is R2 and current flowing through it is I2
And third resistance is R3 and current flowing through it is I3
If second and third resistor have equal value then equal current will flow through them.
R2 = R3 and I2 = I3
I = I1 + I2 + I3 = I1 + 2I2
6 = 3 + 2I2
I2 = 3/2
So R2 = V/I2 = 12/1.5 = 8 ohm
So R3 = R2 = 8 ohm
Hope it clears. If you like answer then please approve it.